# CoMat: Aligning Text-to-Image Diffusion Model with Image-to-Text Concept Matching

### v2

@@ Abstract
Diffusion models have demonstrated great success in the field of text-to-image generation. However, alleviating the misalignment between the text prompts and images is still challenging. We break down the problem into two causes: concept ignorance and concept mismapping. To tackle the two challenges, we propose CoMat, an end-to-end diffusion model fine-tuning strategy with the image-to-text concept matching mechanism. Firstly, we introduce a novel image-to-text concept activation module to guide the diffusion model in revisiting ignored concepts. Additionally, an attribute concentration module is proposed to map the text conditions of each entity to its corresponding image area correctly. Extensive experimental evaluations, conducted across three distinct text-to-image alignment benchmarks, demonstrate the superior efficacy of our proposed method, CoMat-SDXL, over the baseline model, SDXL [38]. We also show that our method enhances general condition utilization capability and generalizes to the long and complex prompt despite not specifically training on it.
Figure 1: Current text-to-image diffusion model still struggles to produce images well-aligned with text prompts, as shown in the generated images of SDXL [38]. Our proposed method CoMat significantly enhances the baseline model on text condition following, demonstrating superior capability in text-image alignment. All the pairs are generated with the same random seed.
@@ 1 Introduction
The area of text-to-image generation has witnessed considerable progress with the introduction of diffusion models [17, 41, 40, 44, 46] recently. These models have demonstrated remarkable performance in creating high-fidelity and diverse images based on textual prompts. However, it still remains challenging for these models to faithfully align with the prompts, especially for the complex ones. For example, as shown in Fig. 1, current state-of-the-art open-sourced model SDXL [38] fails to generate entities or attributes mentioned in the prompts, e.g., the feathers made of lace and dwarfs in the top row. Additionally, it fails to understand the relationship in the prompt. In the middle row of Fig. 1, it mistakenly generates a Victorian gentleman and a quilt with a river on it.
@@ 2 Related Work
Text-to-image alignment is the problem of enhancing coherence between the prompts and the generated images, which involves multiple aspects including existence, attribute binding, relationship, etc. Recent methods address the problem mainly in three ways.
@@
Moreover, some works aim to enhance the alignment using feedback from image understanding models. [22, 48] fine-tune the diffusion model with well-aligned generated images chosen by the VQA model [27] to strategically bias the generation distribution. Other works propose to optimize the diffusion models in an online manner. [13, 4] introduce RL fine-tuning for generic rewards. As for differentiable reward, [11, 59, 57] propose to backpropagate the reward function gradient through the denoising process. Similar to our work, [14] also proposes to leverage image captioning models. We discuss the difference between their method with ours in Appendix C
@@ 4.1 Concept Activation
As noted in Section 1, the diffusion model occasionally exhibits little attention on certain concepts, and the corresponding concept is therefore missing in the image, which we termed as the condition ignorance problem. To address this, our key insight is to add supervision on the generated image to detect the missing concepts. We achieve this by leveraging the image understanding ability of an image-to-text model, which can accurately identify concepts not present in the generated image based on the given text prompt. With the image-to-text model’s supervision, the diffusion model is compelled to revisit text tokens to search for ignored condition information and assign more significance to the previously overlooked text concepts for better text-image alignment. Concretely, given a prompt $\mathcal{P}$ with word tokens $\{w_{1},w_{2},\dots,w_{L}\}$, we first generate an image $\mathcal{I}$ with the denoising function $\epsilon_{\theta}$ after $T$ denoising steps. Then, a frozen image-to-text model $\mathcal{C}$ is used to score the alignment between the prompt and the image in the form of log-likelihood. Therefore, our training objective aims to minimize the negative of the score, denoted as $\mathcal{L}_{i2t}$:
@@
It is important to note that the concepts in the image include a broad field. This method provides a universal solution to various misalignment problems like object existence, complex relationships, etc. To conduct the gradient update through the whole iterative denoising process, we follow [57] to fine-tune the denoising network $\epsilon_{\theta}$, which ensures the training effectiveness and efficiency by simply stopping the gradient of the denoising network input.
@@
Mixed Latent Strategy. Besides, we inject information from real-world images to guide the learning process. Specifically, in addition to the latents starting from pure noise (marked as black in Fig. 4), we obtain the noisy real latents by adding noise on a real-world image at a random timestep $\tau$ (marked as ‘Noisy GT’). We jointly denoise these two types of latents and calculate the loss given by the image-to-text model. The intuition is that, since the noisy real latent is a perturbed version of the real-world image, which is well aligned with its prompt, this provides a shortcut for the diffusion model to directly reconstruct the original image. This guidance can not only smooth the optimization process, but also prohibits the gradient from simply hacking the image-to-text model and encourages the diffusion model to generate an image both aligned with the prompt and of high fidelity.
@@ Appendix A Algorithm
The pseudocode for a single loss computation step for the online T2I-Model is described in  Algorithm 1.
@@ Appendix C More Related Work
The image-to-text model in the main text refers to the models capable of image captioning. Previous image captioning models are pre-trained on various vision and language tasks (e.g., image-text matching, (masked) language modeling) [29, 35, 23, 49], then fine-tuned with image captioning tasks [9]. Various model architectures have been proposed [52, 62, 28, 27, 51]. BLIP [27] takes a fused encoder architecture, while GIT [51] adopts a unified transformer architecture. Recently, multimodel large language models have been flourishing [34, 68, 64, 1]. For example, LLaVA [34] leverages LLM as the text decoder and obtains impressive results. When prompted properly, these models are also able to provide a detailed image caption.
@@
Appendix D Limitation
How to more effectively incorporate Multimodal Large Language Models (MLLMs) into text-to-image diffusion models by our proposed method requires more exploration. MLLM possesses state-of-the-art image-text understanding capability in addition to image captioning. We will focus on leveraging MLLMs to provide finer-grained guidance to the diffusion model in our future work. In addition, we also observe the potential of CoMat for adapting to 3D domains, promoting the text-to-3D generation with stronger alignment. As for the training cost, since our method needs the diffusion model to perform the whole inference process, the training time is extended. Our future direction will be to accelerate the training process.
@@ E.1 Implementation Details
Training Details. In our method, we inject LoRA [19] layers into the UNet of the online training model and discriminator and keep all other components frozen. For both SDXL and SD1.5, we train 2,000 iters on 8 NVIDIA A100 GPUS. We use a local batch size of 6 for SDXL and 4 for SD1.5. We choose Grounded-SAM [43] from other open-vocabulary segmentation models [66, 69]. The DDPM [17] sampler with 50 steps is used to generate the image for both the online training model and the original model. In particular, we follow [57] and only enable gradients in 5 steps out of those 50 steps, where the attribute concentration module would also be operated. Besides, to speed up training, we use training prompts to generate and save the generated latents of the pre-trained model in advance, which are later input to the discriminator during fine-tuning. Training Resolutions. We observe that training SDXL is very slow due to the large memory overhead at $1024\times 1024$. However, SDXL is known to generate low-quality images at resolution $512\times 512$. This largely affects the image understanding of the image-to-text model. So we first equip the training model with better image generation capability at $512\times 512$. We use our training prompts to generate $1024\times 1024$ images with pre-trained SDXL. Then we resize these images to $512\times 512$ and use them to fine-tune the UNet of SDXL for 100 steps, after which the model can already generate high-quality $512\times 512$ images. We continue to implement our method on the fine-tuned UNet.
Training Layers for Attribute Concentration. Only cross-attention maps in the middle blocks and decoder blocks are used to compute the loss.

### v3

@@ Abstract
Diffusion models have demonstrated great success in the field of text-to-image generation. However, alleviating the misalignment between the text prompts and images is still challenging. We break down the problem into two causes: concept ignorance and concept mismapping. To tackle the two challenges, we propose CoMat, an end-to-end diffusion model fine-tuning strategy with the image-to-text concept matching mechanism. Firstly, we introduce a novel image-to-text concept activation module to guide the diffusion model in revisiting ignored concepts. Additionally, an attribute concentration module is proposed to map the text conditions of each entity to its corresponding image area correctly. Extensive experimental evaluations, conducted across three distinct text-to-image alignment benchmarks, demonstrate the superior efficacy of our proposed method, CoMat-SDXL, over the baseline model, SDXL [49]. We also show that our method enhances general condition utilization capability and generalizes to the long and complex prompt despite not specifically training on it. The code is available at https://github.com/CaraJ7/CoMat.
Figure 1: Current text-to-image diffusion model still struggles to produce images well-aligned with text prompts, as shown in the generated images of SDXL [49]. Our proposed method, CoMat, significantly enhances the baseline model on text condition following, demonstrating superior capability in text-image alignment. All the pairs are generated with the same random seed.
@@ 1 Introduction
The area of text-to-image generation has witnessed considerable progress with the introduction of diffusion models [23, 52, 51, 56, 59, 76] recently. These models have demonstrated remarkable performance in creating high-fidelity and diverse images based on textual prompts. However, it still remains challenging for these models to faithfully align with the prompts, especially for the complex ones. For example, as shown in Fig. 1, current state-of-the-art open-sourced model SDXL [49] fails to generate entities or attributes mentioned in the prompts, e.g., the feathers made of lace and dwarfs in the top row. Additionally, it fails to understand the relationship in the prompt. In the middle row of Fig. 1, it mistakenly generates a Victorian gentleman and a quilt with a river on it.
@@ 2 Related Work
Recently, text-to-image diffusion models [56, 52, 65, 66] have become extremely trending, but they have also brought many new challenges [28, 61]. Among them, the text-to-image alignment problem has gained much attention. The problem is defined as the incoherence between the prompts and the generated images, which involves multiple aspects including existence, attribute binding, relationship, etc. Recent methods address the problem mainly in three ways.
@@
Moreover, some works aim to enhance the alignment using feedback from image understanding models. [28, 62] fine-tune the diffusion model with well-aligned generated images chosen by the VQA model [35] to strategically bias the generation distribution. Other works propose to optimize the diffusion models in an online manner. [17, 4] introduce RL fine-tuning for generic rewards. As for differentiable reward, [14, 75, 73] propose to backpropagate the reward function gradient through the denoising process. Other works like [46] enhance the prompt encoding to foster better alignment. Similar to our work, [18] also proposes to leverage image captioning models. We discuss the difference between their method with ours in Appendix C.
@@ 4.1 Concept Activation
As noted in Section 1, the diffusion model occasionally exhibits little attention on certain concepts, and the corresponding concept is therefore missing in the image, which we termed as the condition ignorance problem. To address this, our key insight is to add supervision on the generated image to detect the missing concepts. We achieve this by leveraging the image understanding ability of an image-to-text model, which can accurately identify concepts not present in the generated image based on the given text prompt. With the image-to-text model’s supervision, the diffusion model is compelled to revisit text tokens to search for ignored condition information and assign more significance to the previously overlooked text concepts for better text-image alignment. Concretely, given a prompt $\mathcal{P}$ with word tokens $\{w_{1},w_{2},\dots,w_{L}\}$, we first generate an image $\mathcal{I}$ with the denoising function $\epsilon_{\theta}$ after $T$ denoising steps. Then, a frozen image-to-text model $\mathcal{C}$ is used to score the alignment between the prompt and the image in the form of log-likelihood. The scoring capability of $\mathcal{C}$ comes with the image-to-text models’ training nature. These models are trained for the image captioning task with the negative loglikelihood loss, i.e., the model needs to maximize the probability of generating the caption given the corresponding image. Therefore, whenever the generated image does not align with the text prompt, the model will output a low log-likelihood. Our training objective aims to minimize the negative of the log-likelihood, denoted as $\mathcal{L}_{i2t}$:
@@
Besides, it is also important to note that the concepts in the image include a broad field. This method provides a universal solution to various misalignment problems like object existence, complex relationships, etc. Finally, to conduct the gradient update through the whole iterative denoising process, we follow [73] to fine-tune the denoising network $\epsilon_{\theta}$, which ensures the training effectiveness and efficiency by simply stopping the gradient of the denoising network input.
@@
Mixed Latent Strategy. Besides, we inject information from real-world images to guide the learning process. Specifically, in addition to the latents starting from pure noise (marked as black in Fig. 4), we obtain the noisy real latents by adding noise on a real-world image at a random timestep $\tau$ (marked as ‘Noisy GT’). We jointly denoise these two types of latents and calculate the loss given by the image-to-text model. The intuition is that, since the noisy real latent is a perturbed version of the real-world image, which is well aligned with its prompt, this provides a shortcut for the diffusion model to directly reconstruct the original image. This guidance can not only smooth the optimization process, but also prohibits the gradient from simply hacking the image-to-text model and encourages the diffusion model to generate an image both aligned with the prompt and of high fidelity. More illustration is included in Appendix A.
@@ 5.1 Experimental Setup
StructureDiffusion [19] 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355
Composable Diffusion [44] 0.4063 0.3299 0.3645 0.0800 0.2980 0.2898
Attend-and-Excite [6] 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401
TokenCompose [70] 0.5055 0.4852 0.5881 0.1815 0.3173 0.2937
@@ 5.4 Ablation Study
5.5 Robustness Analysis
We test the robustness of our method by the method proposed in [16], which introduces an automated way to discover prompts that induce misalignment in Stable Diffusion models. We evaluate this attack method on SD1.5 and CoMat-SD1.5 using both short and long prompts.
Table 6: Success rate of prompt attack.
Model Short $\downarrow$ Long $\downarrow$
SD1.5 [56] 49.1% 51.1%
CoMat-SD1.5 46.8% 50.4%
For 1,000 ImageNet-1K classes, we generate 20 samples per class using the attack method and measure the success rate - defined as the proportion of generated images that could be mistakenly classified by a visual classifier. Table 6 shows that CoMat-SD1.5 exhibits lower attack success rates for both prompt lengths, demonstrating enhanced alignment robustness compared to the base model.
6 Limitations
How to more effectively incorporate Multimodal Large Language Models (MLLMs) into text-to-image diffusion models by our proposed method requires more exploration. MLLM possesses state-of-the-art image-text understanding capability in addition to image captioning. We will focus on leveraging MLLMs to provide finer-grained guidance to the diffusion model in our future work. In addition, the attribute concentration module cannot assign attributes to multiple same-name objects, such as an Asian girl with an Indian girl, the segmentation model cannot differentiate two girls and therefore cannot assign attributes. As for the training cost, since our method needs the diffusion model to perform the whole inference process, the training time is extended. Our future direction will be to accelerate the training process.
@@ 7 Conclusion
Acknoledgement
This project is funded in part by National Key R&D Program of China Project 2022ZD0161100, by the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK, by General Research Fund of Hong Kong RGC Project 14204021. Hongsheng Li is a PI of CPII under the InnoHK.
@@ Appendix A Algorithm
Here we first detail the mixed latent training strategy and then provide the pseudocode for a single loss computation step.
The mixed latent strategy contains two types of latents in the fine-tuning procedure, i.e., the latent starting from the pure noise and the noisy latent from the GT Images.
Latent starting from the pure noise This serves as the main branch in our pipeline. Our fine-tuning process shares the same procedure to generate an image as the diffusion model does in the inference time. We uniformly sample $K$ steps from all the inference steps to enable the gradient. Therefore, the latent is sampled from the pure noise $\mathcal{N}(0,I)$. We iteratively denoise it to obtain the generated image. The image is then used to calculate the $\mathcal{L}_{i2t}$ and $\mathcal{L}_{adv}$ loss. It is also sent to the segmentation model to provide the object mask for computing the $\mathcal{L}_{pos}$ and $\mathcal{L}_{neg}$. The latent starting from the pure noise corresponds to the upper left part in Fig. 4. Please refer to [73] for how to receive the gradient from the loss.
Noisy latent from the GT Images We also aim to inject information from the GT images to stabilize the fine-tuning process. We randomly sample a timestamp $\tau$ from a pre-defined range $[T_{1},T_{2}]$. Then we obtain $x_{\tau}$ by adding the timestamped noise $\epsilon_{\tau}$ on the latent of the GT Image $x_{0}$. We also iteratively denoise this noisy GT latent to get $\hat{x}_{0}$ as we do for the latents starting from the pure noise. This $\hat{x}_{0}$ is only used to calculate the $\mathcal{L}_{i2t}$ loss. The latent starting from the noisy GT corresponds to the bottom left part in Fig. 4.
The pseudocode for a single loss computation step for the online T2I-Model is described below.
@@ Appendix C More Related Work
The image-to-text model in the main text refers to the models capable of image captioning. Previous image captioning models are pre-trained on various vision and language tasks (e.g., image-text matching, (masked) language modeling) [37, 45, 31, 64], then fine-tuned with image captioning tasks [12]. Various model architectures have been proposed [68, 78, 36, 35, 67]. BLIP [35] takes a fused encoder architecture, while GIT [67] adopts a unified transformer architecture. Recently, multimodal large language models (MLLMs) have been flourishing [43, 86, 82, 1, 87, 60, 30]. Empowered by the strong language ability of large language models (LLMs) [82, 80], MLLMs are capable of various vision-language tasks like detailed image captioning [43, 86, 63, 8], visual question answering [38, 83, 85, 9, 29], etc. LLaVA [43, 34, 34] is one of the representative MLLMs. When prompted properly, it can generate elaborate image captions.
@@
C.3 vs. Class-specific Prior Preservation Loss [58]
Similarity: Both the class-specific prior preservation loss (CPP Loss) [58] and our proposed fidelity preservation module (FP) share the similar high-level idea of preserving the generation quality while fine-tuning the diffusion models.
Difference: (1) Target task and preserve domain. [58] seeks to personalize image generation for specific objects. While the introduced CPP Loss primarily maintains generative capabilities within a narrow domain—specifically, the object class present in the training data—our proposed FP module operates within the context of text-image alignment. FP aims to preserve general generative capabilities by computing adversarial loss across the entire training dataset, encompassing a diverse range of text prompts. (2) Methodology. Since the training data of [58] finetunes the diffusion model with the pretraining loss, i.e., the squared error denoising loss on a certain timestamp. CPP Loss follows its form. In contrast, our fine-tuning procedure simulates the inference process of the diffusion model to conduct a full-step inference. We aim to directly supervise the generated image to achieve the training-test alignment. Therefore, we propose the novel FP module to leverage a discriminator to adversarially preserve its quality. The applied discriminator is also updated along with the fine-tuning process, enabling finer control of the image quality.
Appendix D Future Work
We believe our work can also be applied in the text-to-video diffusion models. With the introduction of various MLLMs handling videos [81, 10] and video segmentation models [54, 21], both our concept activation and attribute concentration modules could be used for text-video-alignment training.
@@ E.1 Implementation Details
Training Details. In our method, we inject LoRA [25] layers into the UNet of the online training model and discriminator and keep all other components frozen. For both SDXL and SD1.5, we train 2,000 iters on 8 NVIDIA A100 GPUS. We use a local batch size of 6 for SDXL and 4 for SD1.5. We choose Grounded-SAM [55] from other open-vocabulary segmentation models [84, 88]. The DDPM [23] sampler with 50 steps is used to generate the image for both the online training model and the original model. In particular, we follow [73] and only enable gradients in 5 steps out of those 50 steps, where the attribute concentration module would also be operated. Besides, to speed up training, we use training prompts to generate and save the generated latents of the pre-trained model in advance, which are later input to the discriminator during fine-tuning.
Training Resolutions. We observe that training SDXL is very slow due to the large memory overhead at $1024\times 1024$. However, SDXL is known to generate low-quality images at resolution $512\times 512$. This largely affects the image understanding of the image-to-text model. So we first equip the training model with better image generation capability at $512\times 512$. We use our training prompts to generate $1024\times 1024$ images with pre-trained SDXL. Then we resize these images to $512\times 512$ and use them to fine-tune the UNet of SDXL for 100 steps, after which the model can already generate high-quality $512\times 512$ images. We continue to implement our method on the fine-tuned UNet.
Training Layers for Attribute Concentration. Following [70], only cross-attention maps in the middle blocks and decoder blocks are used to compute the loss.